# Recurrent Memory for Online Interdomain Gaussian Processes

### v2

@@ 5 Conclusion
We introduce OHSGPR, a novel online Gaussian process model that leverages the HiPPO framework for robust long-range memory in online regression. By interpreting HiPPO’s time-varying orthogonal projections as interdomain GP basis functions, we establish a link between SSMs and GPs. This connection allows OHSGPR to harness HiPPO’s efficient ODE-based recurrent updates while preserving GP-based uncertainty quantification. Empirical results on time series regression tasks show that OHSGPR significantly outperforms existing online GP methods, especially in scenarios requiring long-term memory. Moreover, with fixed kernel hyperparameters, its recurrence-based kernel updates yield far lower computational overhead than OSGPR’s sequential inducing point optimization. This efficient streaming capability and preservation of historical information make OHSGPR well-suited for real-world applications demanding both speed and accuracy. Future work includes exploring alternative basis and measure selections and integrating OHSGPR with deep learning (e.g., enabling online learning in GPVAE models (Fortuin et al., 2020; Jazbec et al., 2021)).
@@ Appendix A HiPPO-LegS matrices
Here we provide the explicit form of matrices used in our implementation of HiPPO-LegS (Gu et al., 2020). For a given time $t$, the measure $\omega^{(t)}(x)=\frac{1}{t}\mathbf{1}_{[0,t]}(x)$ and basis functions $\phi_{m}^{(t)}(x)=g_{m}^{(t)}(x)\omega^{(t)}(x)=\frac{\sqrt{2m+1}}{t}P_{m}\left(\frac{2x}{t}-1\right)\mathbf{1}_{[0,t]}(x)$ are used, where $P_{m}(\cdot)$ is the $m$-th Legendre polynomial and $\mathbf{1}_{[0,t]}(x)$ is the indicator function on the interval $[0,t]$. These basis functions are orthonormal, i.e.,
@@
Following Gu et al. (2020), the HiPPO-LegS framework maintains a coefficient vector $\mathbf{c}(t)\in\mathbb{R}^{M\times 1}$ that evolves according to the ODE:
@@
$[\mathbf{A(t)}]_{nk}=\begin{cases}-\frac{1}{t}\sqrt{(2n+1)(2k+1)}&\text{if }n>k\\ -\frac{n+1}{t}&\text{if }n=k\\ 0&\text{if }n<k\end{cases}$ (9)
@@
$[\mathbf{B(t)}]_{n}=\frac{\sqrt{2n+1}}{t}$ (10)
@@ Appendix B Computing prior covariance of the inducing variables $\mathbf{K}_{\mathbf{uu}}^{(t)}$
We provide the detailed derivation for the following two approaches when the inducing functions are defined via HiPPO-LegS. Recall that
$[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}=\iint k(x,x^{\prime})\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime},$ (11)
@@ B.1 RFF approximation
For stationary kernels (e.g. an RBF kernel), Bochner’s theorem permits a representation of the kernel as an expectation with respect to its spectral measure:
$k(x,x^{\prime})=\mathbb{E}_{p(w)}\Bigl{[}\cos(wx)\,\cos(wx^{\prime})+\sin(wx)\,\sin(wx^{\prime})\Bigr{]},$ (12)
where $p(w)$ is the spectral density of the kernel.
Substituting this into equation 11, for a given Monte Carlo sample of $w$ we define
$Z_{w,\ell}^{(t)}=\int\cos(wx)\phi_{\ell}^{(t)}(x)\,\mathrm{d}x,\quad Z_{w,\ell}^{\prime(t)}=\int\sin(wx)\phi_{\ell}^{(t)}(x)\,\mathrm{d}x,$ (13)
and
$\mathbf{Z}_{w}^{(t)}=\left[Z_{w,1}^{(t)},\cdots,Z_{w,M}^{(t)}\right]^{\intercal},\quad\mathbf{Z}_{w}^{\prime(t)}=\left[Z_{w,1}^{\prime(t)},\cdots,Z_{w,M}^{\prime(t)}\right]^{\intercal}.$ (14)
Collecting $N$ Monte Carlo samples, we form the feature matrix
@@
$\mathbf{K}_{\mathbf{uu}}^{(t)}\approx\frac{1}{N}\,\mathbf{Z}^{(t)}\left(\mathbf{Z}^{(t)}\right)^{\intercal}.$ (16)
Moreover, as before, the evolution of $\mathbf{Z}_{w}^{(t)}$ and $\mathbf{Z}_{w}^{\prime(t)}$ is governed by the ODE
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{Z}_{w}^{(t)}=\mathbf{A}(t)\,\mathbf{Z}_{w}^{(t)}+\mathbf{B}(t)\,h(t),\quad\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{Z}_{w}^{\prime(t)}=\mathbf{A}(t)\,\mathbf{Z}_{w}^{\prime(t)}+\mathbf{B}(t)\,h^{\prime}(t),$ (17)
with $h(t)=\cos(wt)$ and ${h}^{\prime}(t)=\sin(wt)$ and these ODEs can be solved in parallel for different Monte Carlo samples, thereby fully exploiting the computational advantages of GPU architectures.
@@ B.2 Direct ODE evolution
Differentiating $[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}$ with respect to $t$ gives
$\frac{d}{dt}\,[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}=\iint k(x,x^{\prime})\frac{\mathrm{d}}{\mathrm{d}t}\left[\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\right]\mathrm{d}x\mathrm{d}x^{\prime}.$ (18)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}=\iint k(x,x^{\prime})\frac{\mathrm{d}}{\mathrm{d}t}\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}+\iint k(x,x^{\prime})\phi_{\ell}^{(t)}(x)\frac{\mathrm{d}}{\mathrm{d}t}\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}.$ (19)
In HiPPO-LegS, each $\phi_{\ell}^{(t)}(x)$ obeys an ODE governed by scaled Legendre polynomials on $[0,t]$ and a Dirac boundary term at $x=t$. Concretely,
$\frac{\mathrm{d}}{\mathrm{d}t}\phi_{\ell}^{(t)}(x)=-\frac{\sqrt{2\ell+1}}{t}\left[\frac{\ell+1}{\sqrt{2\ell+1}}\phi_{\ell}^{(t)}(x)+\sqrt{2\ell-1}\phi_{\ell-1}^{(t)}(x)\right]+\frac{1}{t}\,\delta_{t}(x),$ (20)
where $\delta_{t}(x)$ is the Dirac delta at $x=t$.
Substituting this expression into the integrals yields the boundary terms $\int k(t,x^{\prime})\,\phi_{m}^{(t)}(x^{\prime})\,\mathrm{d}x^{\prime}$, along with lower-order terms involving $[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell-1,m}$, etc. Summarizing in matrix form leads to
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathbf{K}_{\mathbf{uu}}^{(t)}=-\frac{1}{t}\left[\mathbf{A}\,\mathbf{K}_{\mathbf{uu}}^{(t)}+\mathbf{K}_{\mathbf{uu}}^{(t)}\mathbf{A}^{\intercal}\right]+\frac{1}{t}\left[\tilde{\mathbf{B}}(t)+\tilde{\mathbf{B}}(t)^{\intercal}\right],$ (21)
where $\mathbf{K}_{\mathbf{uu}}^{(t)}\in\mathbb{R}^{M\times M}$ has entries $[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}$, $\mathbf{A}\in\mathbb{R}^{M\times M}$ is the same lower-triangular matrix from the HiPPO-LegS framework defined in equation equation 9, and $\tilde{\mathbf{B}}(t)\in\mathbb{R}^{M\times M}$ is built from the boundary contributions as
@@
where $\mathbf{c}(t)\in\mathbb{R}^{M\times 1}$ is the coefficient vector and $\mathbf{1}_{M}\in\mathbb{R}^{1\times M}$ is a row vector of ones of size $M$ with
@@ B.2.1 Efficient computation of $\tilde{\mathbf{B}}(t)$
Assuming that the kernel $k(x,t)$ is stationary, it depends only on the difference $d=|x-t|$, so $k(x,t)=k(d)$. In our context, since we integrate over $x\in[t_{\text{start}},t]$ with $x\leq t$, we have $d=t-x\geq 0$. Therefore, we can express $k(x,t)$ as a function of $d$ over the interval $[0,t-t_{\text{start}}]$:
@@
Our goal is to approximate $k(d)$ over the interval $[0,t-t_{\text{start}}]$ using the orthonormal Legendre basis functions scaled to this interval. Specifically, we can represent $k(d)$ as
@@
where $g_{m}^{(t)}(d)$ are the Legendre polynomials rescaled to the interval $[0,t-t_{\text{start}}]$.
@@
To efficiently compute the coefficients $\tilde{c}_{m}(t)$, we utilize the HiPPO-LegS framework, which provides a method for recursively updating the coefficients of a function projected onto an orthogonal basis as the interval expands. In our case, as $t$ increases, the interval $[t_{\text{start}},t]$ over which $k(d)$ is defined also expands, and we can update $\tilde{c}_{m}(t)$ recursively.
@@
where $\tilde{\mathbf{c}}_{k}=\left[\tilde{c}_{0}(t_{k}),\tilde{c}_{1}(t_{k}),\ldots,\tilde{c}_{M-1}(t_{k})\right]^{\intercal}$, and $\mathbf{A}\in\mathbb{R}^{M\times M}$ and $\mathbf{B}\in\mathbb{R}^{M}$ are again matrices defined by the HiPPO-LegS operator (Gu et al., 2020).
@@
The change of variables from $x$ to $d=t-x$ introduces a reflection in the function domain. Since the Legendre polynomials have definite parity, specifically,
@@
This parity correction ensures that the computed coefficients properly represent the function over the interval $[t_{\text{start}},t]$ without the effect of the reflection.
@@
Appendix C Performance of OHSGPR and baselines with fixed kernel
In Figure 3 and 4, we present the performance, measured by RMSE and NLPD, of OSGPR, OVFF and OHSGPR models with fixed kernel shown in Section 4.2 for Solar Irradiance and Audio dataset, respectively. OHSGPR again outperforms OSGPR in terms of long-term memory preservation and achieves better RMSE and NLPD at the later stage where most of the tasks have been revealed. Interestingly, the overall performance of these fixed-kernel models is competitive or even better with the models with trainable kernel in Section 4. The performance of OVFF with fixed kernel is significantly better than OVFF with trainable kernel, suggesting it may be hard to find good hyperparamters with the online ELBO objective for OVFF. Furthermore, compared to OHSGPR with a trainable kernel, OHSGPR with a fixed kernel achieves better performance on the audio dataset when we increase the total number of discretizations from 160 to 320 or 480. The best RMSE and NLPD for OHSGPR after the final task for audio data in this work (0.2738 and 0.5510, respectively) are achieved by a fixed-kernel OHSGPR with 480 total discretization steps (Figure 4(d) and 4(h)), suggesting in practice, for some applications, if the kernel hyperparameters are already in a good range, further optimization over it might show diminishing returns and in this case the effect of fine-grained discretization might play a more critical role in the performance.

### v4

@@ 6 Conclusion
We introduce OHSVGP, a novel online Gaussian process model that leverages the HiPPO framework for robust long-range memory in online/continual learning. By interpreting HiPPO’s time-varying orthogonal projections as adaptive interdomain GP basis functions, we leverage SSM for improved online GP. This connection allows OHSVGP to harness HiPPO’s efficient ODE-based recurrent updates while preserving GP-based uncertainty-aware prediction. Empirical results on a suite of online and continual learning tasks show that OHSVGP outperforms existing online sparse GP methods, especially in scenarios requiring long-term memory. Moreover, its recurrence-based covariance updates yield far lower computational overhead than OSVGP’s sequential inducing point optimization. This efficient streaming capability and preservation of historical information make OHSVGP well-suited for real-world applications demanding both speed and accuracy.
Broader impact.
This paper presents work whose goal is to advance machine learning research. There may exist potential societal consequences of our work, however, none of which we feel must be specifically highlighted here.
Acknowledgments and Disclosure of Funding
Samir Bhatt acknowledges funding from the MRC Centre for Global Infectious Disease Analysis (reference MR/X020258/1), funded by the UK Medical Research Council (MRC). This UK funded award is carried out in the frame of the Global Health EDCTP3 Joint Undertaking. Samir Bhatt acknowledges support from the Danish National Research Foundation via a chair grant (DNRF160) which also supports Jacob Curran-Sebastian. Samir Bhatt acknowledges support from The Eric and Wendy Schmidt Fund For Strategic Innovation via the Schmidt Polymath Award (G-22-63345) which also supports Harrison Bo Hua Zhu. Samir Bhatt acknowledges support from the Novo Nordisk Foundation via The Novo Nordisk Young Investigator Award (NNF20OC0059309).
@@ Appendix A HiPPO-LegS matrices
Here we provide the explicit form of matrices used in our implementation of HiPPO-LegS [Gu et al., 2020]. For a given time $t$, the measure $\omega^{(t)}(x)=\frac{1}{t}\mathbf{1}_{[0,t]}(x)$ and basis functions $\phi_{m}^{(t)}(x)=g_{m}^{(t)}(x)\omega^{(t)}(x)=\frac{\sqrt{2m+1}}{t}P_{m}\left(\frac{2x}{t}-1\right)\mathbf{1}_{[0,t]}(x)$ are used, where $P_{m}(\cdot)$ is the $m$-th Legendre polynomial and $\mathbf{1}_{[0,t]}(x)$ is the indicator function on the interval $[0,t]$. These basis functions are orthonormal, i.e.,
@@
Following Gu et al. [2020], the HiPPO-LegS framework maintains a coefficient vector $\mathbf{c}(t)\in\mathbb{R}^{M\times 1}$ that evolves according to the ODE:
@@
[…]
@@
$\mathbf{B}(t)]_{n}=\frac{[\mathbf{B}]_{n}}{t}=\frac{\sqrt{2n+1}}{t}$ (9)
@@ Appendix B Computing prior covariance of the inducing variables $\mathbf{K}_{\mathbf{uu}}^{(t)}$
We provide the detailed derivation for the following two approaches when the inducing functions are defined via HiPPO-LegS. Recall that the $\ell m$-th element of the prior covariance matrix for the inducing variables is given by
$\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}=\iint k(x,x^{\prime})\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime},$ (10)
@@ B.1 RFF approximation
Since $k(x,x^{\prime})$ depends on both $x$ and $x^{\prime}$, a recurrence update rule based on the original HiPPO formulation, which is designed for single integral, can not be obtained directly for $\bm{K}_{\bm{uu}}^{(t)}$. Fortunately, for stationary kernels, Bochner Theorem [Rudin, 1994] can be applied to factorize the above double integrals into two separate single integrals, which gives rise to Random Fourier Features (RFF) approximation [Rahimi and Recht, 2007]: for a stationary kernel $k(x,x^{\prime})=k(|x-x^{\prime}|)$, RFF approximates it as follows:
$\begin{split}k(x,x^{\prime})&=\mathbb{E}_{p(w)}\Bigl[\cos(wx)\,\cos(wx^{\prime})+\sin(wx)\,\sin(wx^{\prime})\Bigr]\\ &\approx\frac{1}{N}\sum_{n=1}^{N}\left[\cos\left(w_{n}x\right)\cos\left(w_{n}x^{\prime}\right)+\sin\left(w_{n}x\right)\sin\left(w_{n}x^{\prime}\right)\right],\end{split}$ (11)
where $w_{n}\sim p(w)$ is the spectral density of the kernel. Substituting this into the double integral (Eq. 10) factorizes the dependency on $x$ and $x^{\prime}$, reducing $\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell m}$ to addition of products of one-dimensional integrals. Each integral based on a Monte Carlo sample $w$ has the form of either
$Z_{w,\ell}^{(t)}=\int\cos(wx)\phi_{\ell}^{(t)}(x)\,\mathrm{d}x\quad\text{or}\quad Z_{w,\ell}^{\prime(t)}=\int\sin(wx)\phi_{\ell}^{(t)}(x)\,\mathrm{d}x,$ (12)
which corresponds to a standard projection coefficient in the HiPPO framework. We further stack these integrals based on $M$ basis functions and define
$\mathbf{Z}_{w}^{(t)}=\left[Z_{w,1}^{(t)},\cdots,Z_{w,M}^{(t)}\right]^{\top},\quad\mathbf{Z}_{w}^{\prime(t)}=\left[Z_{w,1}^{\prime(t)},\cdots,Z_{w,M}^{\prime(t)}\right]^{\top}.$ (13)
Collecting $N$ Monte Carlo samples $\{w_{n}\}_{n=1}^{N}$, we form the feature matrix
@@
$\mathbf{K}_{\mathbf{uu}}^{(t)}\approx\frac{1}{N}\,\mathbf{Z}^{(t)}\left(\mathbf{Z}^{(t)}\right)^{\top}.$ (15)
Since $\mathbf{Z}_{w_{n}}^{(t)}$ and $\mathbf{Z}_{w_{n}}^{\prime(t)}$ are standard HiPPO projection coefficient, their computation is governed by the HiPPO ODE evolution as before
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{Z}_{w_{n}}^{(t)}=\mathbf{A}(t)\,\mathbf{Z}_{w_{n}}^{(t)}+\mathbf{B}(t)\,h_{n}(t),\quad\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{Z}_{w_{n}}^{\prime(t)}=\mathbf{A}(t)\,\mathbf{Z}_{w_{n}}^{\prime(t)}+\mathbf{B}(t)\,h^{\prime}_{n}(t),$ (16)
with $h_{n}(t)=\cos(w_{n}t)$ and ${h}^{\prime}_{n}(t)=\sin(w_{n}t)$ and these ODEs can be solved in parallel across different Monte Carlo samples. In summary, the procedure involves sampling multiple random features, updating them recurrently to time $t$, and averaging across samples to obtain RFF approximation of $\mathbf{K_{uu}}^{(t)}$.
For non-stationary kernels, more advanced Fourier feature approximation techniques (e.g., [Ton et al., 2018]) can be applied.
@@ B.2 Direct ODE evolution
Differentiating $\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell,m}$ with respect to $t$ gives
$\frac{d}{dt}\,[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell,m}=\iint k(x,x^{\prime})\frac{\partial}{\partial t}\left[\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\right]\mathrm{d}x\mathrm{d}x^{\prime}.$ (17)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell,m}=\iint k(x,x^{\prime})\frac{\partial}{\partial t}\phi_{\ell}^{(t)}(x)\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}+\iint k(x,x^{\prime})\phi_{\ell}^{(t)}(x)\frac{\partial}{\partial t}\phi_{m}^{(t)}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}.$ (18)
In HiPPO-LegS, each $\phi_{\ell}^{(t)}(x)$ obeys an ODE governed by lower-order scaled Legendre polynomials on $[0,t]$ and a Dirac delta boundary term at $x=t$. Concretely,
$\begin{split}\frac{\partial}{\partial t}\phi_{\ell}^{(t)}(x)&=-\frac{\sqrt{2\ell+1}}{t}{\LARGE[}\frac{\ell+1}{\sqrt{2\ell+1}}\phi_{\ell}^{(t)}(x)+\sqrt{2\ell-1}\phi_{\ell-1}^{(t)}(x)\\ &\qquad+\sqrt{2\ell-3}\phi_{\ell-2}^{(t)}(x)\cdots{\Large]}+\frac{1}{t}\,\delta_{t}(x),\end{split}$ (19)
where $\delta_{t}(x)$ is the Dirac delta at $x=t$ (see Appendix D.3 in Gu et al. [2020] for details).
Substituting this expression into the integrals yields the boundary terms of the form $\int k(t,x^{\prime})\,\phi_{m}^{(t)}(x^{\prime})\,\mathrm{d}x^{\prime}$, along with lower-order terms involving $\{[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell,m},[\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell-1,m},\cdots\}$, etc. Summarizing in matrix form leads to
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathbf{K}_{\mathbf{uu}}^{(t)}=\left[\mathbf{A}(t)\,\mathbf{K}_{\mathbf{uu}}^{(t)}+\mathbf{K}_{\mathbf{uu}}^{(t)}\mathbf{A}(t)^{\intercal}\right]+\frac{1}{t}\left[\tilde{\mathbf{B}}(t)+\tilde{\mathbf{B}}(t)^{\intercal}\right],$ (20)
where the $lm$-th entry of $\mathbf{K}_{\mathbf{uu}}^{(t)}\in\mathbb{R}^{M\times M}$ is $\mathbf{K}_{\mathbf{uu}}^{(t)}]_{\ell,m}$, $\mathbf{A}(t)\in\mathbb{R}^{M\times M}$ is the same lower-triangular matrix from the HiPPO-LegS framework defined in Eq. 8, and $\tilde{\mathbf{B}}(t)\in\mathbb{R}^{M\times M}$ is built from the boundary contributions as
@@
where $\mathbf{1}_{M}\in\mathbb{R}^{1\times M}$ is a row vector of ones of size $M$ and $\mathbf{c}(t)\in\mathbb{R}^{M\times 1}$ is the coefficient vector with each element being
@@ B.2.1 Efficient computation of $\tilde{\mathbf{B}}(t)$
Assuming that the kernel $k(x,t)$ is stationary, it depends only on the difference d=|xt|d=|x-t|, so $k(x,t)=k(d)$. In our context, since we integrate over $x\in[t_{\text{start}},t]$ with $x\leq t$, we have $d=t-x\geq 0$. Therefore, we can express $k(x,t)$ as a function of $d$ over the interval $0,t-t_{\text{start}}]$:
@@
Our goal is to approximate $k(d)$ over the interval $0,t-t_{\text{start}}]$ using the orthonormal Legendre basis functions scaled to this interval. Specifically, we can represent $k(d)$ as
@@
where $g_{m}^{(t)}(d)$ are the Legendre polynomials rescaled to the interval $0,t-t_{\text{start}}]$.
@@
To efficiently compute the coefficients $\tilde{c}_{m}(t)$, we utilize the HiPPO-LegS framework, which provides a method for recursively updating the coefficients of a function projected onto an orthogonal basis as the interval expands. In our case, as $t$ increases, the interval $t_{\text{start}},t]$ over which $k(d)$ is defined also expands, and we can update $\tilde{c}_{m}(t)$ recursively.
@@
where $\tilde{\mathbf{c}}_{k}=\left[\tilde{c}_{0}(t_{k}),\tilde{c}_{1}(t_{k}),\ldots,\tilde{c}_{M-1}(t_{k})\right]^{\intercal}$, and $\mathbf{A}\in\mathbb{R}^{M\times M}$ and $\mathbf{B}\in\mathbb{R}^{M}$ are again matrices defined by the HiPPO-LegS operator as in Eq. 8 and 9.
@@
The change of variables from $x$ to d=txd=t-x introduces a reflection in the function domain. Since the Legendre polynomials have definite parity, specifically,
@@
This parity correction ensures that the computed coefficients properly represent the function over the interval $t_{\text{start}},t]$ without the effect of the reflection.
@@
B.2.2 Unstability of directly evolving $\mathbf{K}_{\mathbf{uu}}^{(t)}$ as ODE.
Empirically, we find that the direct ODE approach is less stable compared with RFF approach. Intuitively, it can be seen from the difference in the forms of their evolutions, especially in the first term. In RFF approach, the first term of the evolution of Fourier feature is of the form $\mathbf{A}(t)\,\mathbf{Z}_{w}^{(t)}$, which includes evolving vectors with the operator $\mathcal{L}_{1}:\mathbf{X}\rightarrow\mathbf{A}(t)\mathbf{X}$. In direct ODE approach, the first term in the direct evolution of $\mathbf{K}_{\mathbf{uu}}^{(t)}$ is of the form $\mathbf{A}(t)\,\mathbf{K}_{\mathbf{uu}}^{(t)}+\mathbf{K}_{\mathbf{uu}}^{(t)}\mathbf{A}(t)^{\top}$, which requires the Lyapunov operator $\mathcal{L}_{2}:\mathbf{X}\rightarrow\mathbf{A}(t)\mathbf{X}+\mathbf{XA}(t)^{\top}$. The critical difference is that $\mathcal{L}_{2}$ has eigenvalues $\{\lambda_{i}+\lambda_{j}\}$ (where $\lambda_{i}$ and $\lambda_{j}$ are eigenvalues of $\mathbf{A}(t)$) [Horn and Johnson, 1991], while $\mathcal{L}_{1}$ has eigenvalues $\{\lambda_{i}\}$. Since HiPPO-LegS uses a lower-triangular $\mathbf{A}(t)$ with negative diagonal entries, the eigenvalues are all negative $\lambda_{i}<0$. Hence, the eigenvalues of the Lyapunov operator $\mathcal{L}_{2}$ are approximately as twice negative as the eigenvalues of $\mathcal{L}_{1}$, leading to a stiff ODE system with poorer numerical conditioning.
Appendix C Finite basis approximation of posterior OHSVGP
Here, we show that $q(\mathbf{u}^{(t)})$ is the distribution of HiPPO coefficients of the posterior OHSVGP $q_{t}(f)$. From Eq. 2, the posterior of the function values evaluated at arbitrary indices $\mathbf{X}$ is $q_{t}(\mathbf{f_{X}})=\mathcal{N}(\mathbf{f_{X}};\mathbf{K}^{(t)}_{\mathbf{f_{X}}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{(t)-1}\mathbf{m}^{(t)}_{\mathbf{u}},\mathbf{K}_{\mathbf{f_{X}}\mathbf{f_{X}}}-\mathbf{K}^{(t)}_{\mathbf{f_{X}}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{(t)-1}[\mathbf{K}^{(t)}_{\mathbf{u}\mathbf{u}}-\mathbf{S}^{(t)}_{\mathbf{u}}]\mathbf{K}_{\mathbf{u}\mathbf{u}}^{(t)-1}\mathbf{K}^{(t)}_{\mathbf{u}\mathbf{f_{X}}})$.
Based on this, we compute the mean of the $m$-th HiPPO coefficient for $q_{t}(f)$ as follows,